# Generalized Euler Equation from Effective Action: Implications for the Smarr Formula in AdS Black Holes

### v1

@@ Abstract
We derived a generalized Euler equation, $\epsilon+p=sT+\mu q+y\frac{\partial p}{\partial y}$, using the effective field theory formulation of perfect fluids. This generalization was achieved by introducing a new variable $y$ into the effective action, which encodes geometrical information about the spacetime where the fluid is on. Notably, the generalized Euler equation is independent of the AdS/CFT correspondence. However, when applied to a holographic perfect fluid, this equation naturally recovers the Smarr formula for AdS black holes, thus situating the physical interpretation of the Smarr formula within the framework of Gauge/Gravity duality. Finally, our findings raise questions regarding the validity of the extended black hole thermodynamics proposal, which associates the cosmological constant $\Lambda$ with a thermodynamic variable.
@@ 1 Introduction
The AdS/CFT correspondence Juan1 has greatly advanced our understanding of quantum field theory in the strong coupling regime, in part by geometrizing abstract concepts from quantum physics. For instance, a thermal state in the field theory is dual to an eternal AdS black hole Juan2 . Moreover, the Hawking-Page transition between thermal AdS spacetime and a "large" AdS black hole Hawking is interpreted as the confinement/deconfinement transition in a gauge theory on $S^{1}\times S^{3}$ Witten . Within this framework, Karch derived the Smarr formula for AdS black holes from a holographic perspective.
In this article, we revisit the physical meaning of the Smarr formula within the context of gauge/gravity duality. Our goal is to explore the Smarr formula from the perspective of fluid/gravity duality Hubeny1 ; Ranga ; Hubeny2 . It is well-established that the boundary stress tensor of an AdS black hole takes the form of a perfect fluid in global equilibrium Clifford0 ; Clifford . From this, we observe that the thermodynamics of the holographic fluid is, in essence, the thermodynamics of black holes. This opens a bridge between the Smarr formula and the theory of perfect fluids that will investigate.
Another important development relevant to this work is the recent construction of effective actions for perfect fluids Dubovsky1 ; Son ; Dubovsky2 ; Liu2 ; Liu3 ; Ranga1 ; Ranga2 ; Ranga3 ; Ranga4 ; Jensen1 . In this framework, the effective action simultaneously reproduces the perfect fluid stress tensor and ensures the validity of the Euler equation, as the effective Lagrangian functions as a local thermodynamic potential. As we will discuss in Section 2, when the Smarr formula is expressed in a suitable form, it resembles the Euler equation with an additional correction. We interpret this as an indication of the existence of a generalization of the Euler equation, insight that we confirm by deriving a generalized Euler equation from the effective action for perfect fluids. This is achieved by introducing a new variable into the local thermodynamic potential, which is both diffeomorphism-invariant with respect to spacetime transformations and gauge-invariant under the fluid’s emergent low-energy gauge transformations. As we will demonstrate, this new variable has a geometrical interpretation and leads to the central result of this article:
@@
Another context in which the Smarr formula has been extensively discussed is extended black hole thermodynamics (EBHT). In essence, this framework offers a specific interpretation of the Smarr formula by introducing a thermodynamic variable associated with the cosmological constant $\Lambda$ Kastor ; Gibbons ; Dolan1 ; Dolan2 ; Dolan3 . Initially, this new thermodynamic variable was identified as a pressure term $p_{\Lambda}\sim\frac{1}{l^{2}}$, but in more recent versions of EBHT, it has been reinterpreted as the central charge $C\sim N^{2}$ where $N$ represents the number of degree of freedom of the dual field theory Visser1 ; Gao ; Visser2 ; Mann6 . This revision was motivated by the work in Karch , which sought to make EBHT consistent with holography. However, the critical point for this article is that our derivation of the generalized Euler equation—and its physical interpretation—conflicts with the interpretation of the Smarr formula as proposed by EBHT. More detailed comments on this issue will be provided in Section 5.
This paper is organized as follows: In Section 2, we review the Smarr formula and highlight its preliminary similarities with the Euler equation. In Section 3, we examine the standard construction of the effective action for perfect fluids and introduce a new variable into the effective action, which enables us to derive a generalized Euler equation. In Section 4, we demonstrate that the Smarr formula is a specific instance of the generalized Euler equation. Section 5 discusses why our results contradict the proposal for extended black hole thermodynamics. Finally, in Section 6, we present our conclusions and suggest potential directions for future research.
@@ 2 Smarr formula for AdS black holes
where $G_{N}$ is the $d$-dimensional Newton’s constant. The negative cosmological constant $\Lambda$ can be expressed in terms of the AdS radius $l$ as follows: $\Lambda=-\frac{(d-2)(d-1)}{2l^{2}}$. For the spherical AdSd Reissner-Nordström black hole, the metric and gauge field are, respectively,
@@
The constant $\mu$ is eventually identified with the chemical potential, and it is defined such that the gauge field $A$ vanishes at the event horizon located at $r_{h}$
@@
The traditional way of deriving the Smarr formula is by applying Euler’s theorem for homogeneous functions with respect to the "bulk" scaling of thermodynamic quantities [^1] Scaling arguments have been extensively employed in the black hole thermodynamics literature to derive Smarr formulas, accommodating a range of matter contents and asymptotic behaviors (see, e.g., Lu ; Alfaro1 ). A more formal approach to obtaining the Smarr formula involves evaluating the Komar integral at spatial infinity and the horizon. For recent developments in this approach, see Tomas1 ; Tomas2 ; Tomas3 .. For the purposes of this article, it is better to derive the Smarr formula in a different way. Let’s consider the grand canonical potential $\Omega$:
@@
In our context, $M$ represents the internal energy $U$ of the system. It is important to note that, in principle, $\Omega$ is a function of the volume $V$, which, in the case of black holes in AdS spacetimes, corresponds to the spatial volume of the boundary metric where the dual field theory resides. In this work, we interpret black hole thermodynamics as the thermodynamics of the dual field theory in the strongly interacting regime, in line with the AdS/CFT correspondence. For the AdSd Reissner-Nordström black hole, we compute the grand canonical potential $\Omega$ using the Euclidean on-shell action
@@
Equating the definition of $\Omega$ in (9) with the "on-shell value" of $\Omega$ (10), we obtain the Smarr formula for AdS black holes Karch :
@@
In this expression, the event horizon radius $r_{h}$ should be understood as a function of the thermodynamic variables. In summary, the Smarr formula is an unusual relation among thermodynamic quantities.
In this work, we aim to develop a deeper physical understanding of the Smarr formula within the broader framework of the AdS/CFT correspondence. We will demonstrate that the Smarr formula is a natural generalization of the Euler equation, $U+pV=TS+\mu Q$, with an additional correction due to a geometrical scale inherent to the physical problem. While a similar idea is explored in Karch , our approach extends beyond pure thermodynamic arguments by leveraging insights from fluid/gravity duality.
In thermodynamics, the Euler equation arises in the thermodynamic limit, where the volume $V$ and the number of particles $N$ both tend to infinity while maintaining a fixed ratio $\frac{N}{V}$. In this limit, the grand canonical potential becomes $\Omega=-pV$, and the Euler equation follows as $-pV=U-TS-\mu Q$. We seek to extend this reasoning to the Smarr formula. However, in the context of black hole thermodynamics, defining the thermodynamic volume $V$ and pressure $p$ is far from straightforward. This is where fluid/gravity duality offers crucial insight. To define the pressure $p$ for AdS black holes, we observe that the boundary stress-energy tensor associated with these black holes takes the form of a perfect fluid stress tensor Clifford0 ; Clifford . The boundary stress tensor for any given gravitational solution can be calculated using the formula provided in Kraus :
$T_{ab}=\frac{1}{8\pi G_{N}}\left(h_{ab}\left(L-\frac{(d-3)}{l}\right)+K_{ab}+\frac{lG_{ab}}{(d-4)}\right)~{},$ (12)
where $h_{ab}$ is the induced metric, $K_{ab}$ is the extrinsic curvature, and $G_{ab}$ is the Einstein tensor for a fixed value of the radial direction $r$. As a reminder, $d$ is the bulk dimension, so $(d-1)$ is the boundary dimension. In our case, the boundary metric is
@@
Thus, the volume of the spatial section of the boundary spacetime is $V=l^{d-2}A_{d-2}$ where the AdS radius $l$ is acting as the radius of the $(d-2)$-dimensional sphere. The components of the boundary stress tensor are
$T^{t}_{\;t}=\frac{(d-2)\tilde{m}}{16\pi G_{n}l^{2}}=\frac{M}{l^{d-2}A_{d-2}}=\frac{M}{V}~{},$ (14)
$T^{\phi_{j}}_{\;\phi_{j}}=\frac{\tilde{m}}{16\pi G_{n}l^{2}}=\frac{M}{(d-2)l^{d-2}A_{d-2}}=\frac{M}{(d-2)V}~{}.$ (15)
@@
$\epsilon=\frac{M}{V}~{},\quad p=\frac{\epsilon}{(d-2)}~{}.$ (17)
This holographic perfect fluid in certain sense is dual to the AdSd Reissner-Nordström black hole. Besides, this fluid satisfies the conformal equation of state $\epsilon=(d-2)p$. The thermodynamics of this fluid is the same than the black hole thermodynamics since the fluid is at global thermal equilibrium.
@@
which resembles the Euler equation plus an additional correction term.
As mentioned earlier, the traditional Euler equation arises in the thermodynamic limit, where $V\to\infty$. However, in our case, the volume $V$ occupied by the fluid is finite, and thus the standard derivation of the Euler equation no longer applies. In the analysis leading to equation (18), we relied on the idea that a perfect fluid in global thermal equilibrium provides an equivalent description of global thermodynamics. Yet, perfect fluids are generally used to describe local thermodynamics, where the Euler equation is assumed to always hold—a point that appears contradictory to our intuitive understanding of the Smarr formula that we want to verify.
To fully comprehend the physical significance of the Smarr formula, it is necessary to generalize the Euler equation to apply in a broader setting. For this reason, in the next section, we will set aside the discussion of black holes and focus exclusively on the Euler equation in the context of perfect fluids.
@@ 3 Effective action for perfect fluids
In this section, we begin by reviewing the standard construction of the effective action for perfect fluids, which naturally leads to the Euler equation. We then extend this framework by introducing a new variable into the effective action, capturing the fluid’s response to the presence of a geometrical scale. This scale is relevant in certain spacetimes where the fluid resides. With this extension, we derive a generalized form of the Euler equation.
@@ 3.1 Review standard effective action for perfect fluids
Currently, there are formal constructions of effective actions for hydrodynamics using techniques such as the Schwinger-Keldysh formalism Liu2 ; Liu3 ; Ranga1 ; Ranga2 and superspace methods Ranga3 ; Ranga4 ; Jensen1 , with comprehensive reviews available in Liu4 ; Basar . Although these advanced techniques will not be explicitly employed in this article, they highlight the significant progress and depth this field has achieved. Instead, we revisit the foundational constructions presented in Dubovsky1 ; Son ; Dubovsky2 , informed by a more modern understanding of the subject. Moreover, we focus exclusively on perfect fluids, setting aside considerations of dissipation and hydrodynamic anomalies.
To construct the effective action for perfect fluids, two components are required:
@@
Given the dynamical variables and symmetries of the effective action, we should, in principle, consider all gauge-invariant terms in the theory. However, we will disregard higher-derivative terms in the effective action since our interest is limited to perfect fluids.
In the introduction, we assumed that the bulk spacetime has $d$ dimensions. Consequently, in this section, we will consider a perfect fluid residing in a $(d-1)$-dimensional curved spacetime, though our focus will be on the field-theoretical aspects of the discussion. The connection between the results of this section and gravity will be explored in the following section. Let’s begin by establishing the notation: the background metric for the perfect fluid is
@@
These symmetries transformation can be motivated by noticing that $\sigma^{I}$ represents a spatial coordinates describing the fluids. Thus, $\sigma^{I}$ must satisfy the comoving condition
@@
where $x^{a}$ in the position in the spacetime, and the fluid velocity is $u^{a}=\frac{dx^{a}}{d\tau}$. In the second expression of the comoving condition, we use the directional derivative along the fluid velocity $u^{a}$. The comoving condition (23) is preserved by the symmetry transformations (21), (22). It is also important to note that the analogous expression for the temporal scalar field, $\frac{d}{d\tau}\sigma^{0}(x^{a}(\tau))$ is non-zero, but remains gauge-invariant under these symmetry transformations (21), (22).
In essence, the perfect fluid is described by a target metric $G_{MN}(\sigma^{M})$ with fluid coordinates $\sigma^{M}=(\sigma^{0},\sigma^{I})$. In this way, the internal symmetries (21), (22) are the isometries of degenerate metric Son
$ds^{2}=G_{MN}d\sigma^{M}d\sigma^{N}~{},\quad G_{NM}n^{M}=0~{},\quad M,N\in\{0,1,...,d-1\}~{},$ (24)
@@
$ds^{2}=0\cdot(d\sigma^{0})^{2}+G_{IJ}(d\sigma^{I}-v^{I}d\sigma^{0})(d\sigma^{J}-v^{J}d\sigma^{0})~{},\quad I,J\in\{1,...,d-1\}~{},$ (25)
@@
Given the low-energy dynamical variables $\sigma^{a}$, $\varphi$, and the associated low-energy gauge symmetries, we can now construct the effective action for a charged perfect fluid. Using the auxiliary fluid variables $\sigma^{a}$ and $\varphi$, along with the fluid velocity $u^{a}$ and the target metric $G_{MN}$, we form diffeomorphism-invariant scalars with respect to the physical metric $g_{ab}$. These scalars must also be gauge-invariant under the fluid symmetry transformations (21) and (22). The first of these scalars, the entropy density $s$ of the system, is defined in terms of the spatial target metric $G_{IJ}$ (25) as follows Son :
@@
Thus, the Euler equation is an unavoidable consequence of the effective action (32). Note that the equation of state for the fluid is not used in this derivation.
For completeness, we mention that due to the comoving condition, the entropy current $s^{a}$ is conserved off-shell:

### v2

@@ Abstract
We derive a generalized Euler equation, $\epsilon+p=sT+\mu q+y\frac{\partial p}{\partial y}$, using the effective field theory formulation of perfect fluids. This generalization was achieved by introducing a new variable $y$ into the effective action, which encodes a geometrical scale of the spacetime where the fluid is on. Notably, the generalized Euler equation is independent of the AdS/CFT correspondence. However, when applied to a holographic perfect fluid, this equation naturally recovers the Smarr formula for AdS black holes, thus situating the physical interpretation of the Smarr formula within the framework of well-established physics. Finally, our findings raise important questions regarding the validity of treating the cosmological constant $\Lambda$ as a thermodynamic variable, as proposed in certain frameworks within the literature.
@@ 1 Introduction
The AdS/CFT correspondence Juan1 has greatly advanced our understanding of quantum field theory in the strong coupling regime, in part by geometrizing abstract concepts from quantum physics in gravitational terms. For instance, a thermal state in field theory is dual to an eternal AdS black hole Juan2 . Moreover, the Hawking-Page transition between thermal AdS spacetime and a "large" AdS black hole Hawking is interpreted as the confinement/deconfinement transition in a gauge theory on $S^{1}\times S^{3}$ Witten . Within this holographic framework, Karch derived the Smarr formula for AdS black holes, providing further insight into its thermodynamic interpretation.
In this article, we revisit the physical meaning of the Smarr formula within the context of gauge/gravity duality. Our goal is to explore the Smarr formula from the perspective of fluid/gravity duality Hubeny1 ; Ranga ; Hubeny2 . It is well-established that the boundary stress tensor of many AdS black holes takes the form of a perfect fluid in global equilibrium Clifford0 ; Clifford . From this perspective, the thermodynamics of the holographic fluid directly corresponds to black hole thermodynamics. This connection establishes a natural bridge between the Smarr formula and the theory of perfect fluids, which we explore in detail.
A significant development relevant to this work is the recent construction of effective actions for perfect fluids Dubovsky1 ; Son ; Dubovsky2 ; Liu2 ; Liu3 ; Ranga1 ; Ranga2 ; Ranga3 ; Ranga4 ; Jensen1 . Within this framework, the effective action not only reproduces the perfect fluid stress tensor but also ensures the validity of the Euler equation, with the effective Lagrangian acting as a local thermodynamic potential. As we will explore in Section 2, the Smarr formula, when suitably reformulated, bears a close resemblance to the Euler equation—with an additional correction term. This resemblance motivates the pursuit of a generalized Euler equation, which we establish via an explicit derivation from the effective action for perfect fluids. Crucially, this derivation introduces a new variable into the local thermodynamic potential—one that is invariant under both spacetime diffeomorphisms and the emergent low-energy gauge symmetries of the fluid. As we will show, this variable admits a natural geometric interpretation and leads directly to the central result of this work:
@@
Another prominent context in which the Smarr formula has been extensively discussed is extended black hole thermodynamics (EBHT). This framework provides a specific interpretation of the Smarr formula by introducing a thermodynamic variable associated with the cosmological constant $\Lambda$ Kastor ; Gibbons ; Dolan1 ; Dolan2 ; Dolan3 . Initially, this variable was identified as a pressure term, $p_{\Lambda}\sim\frac{1}{l^{2}}$. However, more recent formulations of EBHT reinterpret it as the central charge, $C\sim N^{2}$, where $N$ represents the number of degrees of freedom in the dual field theory Visser1 ; Gao ; Visser2 ; Mann6 . This shift was primarily motivated by Karch , which aimed to align EBHT with holography. However, the central result of this article, (1), challenges these interpretations of the Smarr formula, as we shall discuss later.
This paper is organized as follows: In Section 2, we review the Smarr formula and highlight its preliminary similarities with the Euler equation. In Section 3, we examine the standard construction of the effective action for perfect fluids and introduce a new variable into the effective action, which enables us to derive a generalized Euler equation. In Section 4, we demonstrate that the Smarr formula is a specific instance of the generalized Euler equation. Section 5 discusses why our results are inconsistent with the proposal for extended black hole thermodynamics. Finally, in Section 6, we present our conclusions and suggest potential directions for future research.
@@ 2 Smarr formula for AdS black holes
where $G_{N}$ is the $d$-dimensional Newton’s constant. The negative cosmological constant $\Lambda$ can be expressed in terms of the AdS radius $l$ as follows: $\Lambda=-\frac{(d-2)(d-1)}{2l^{2}}$. For the spherical AdSd Reissner-Nordström black hole with $d>3$, the metric and gauge field are, respectively,
@@
The constant $\mu$ is the chemical potential, and it is defined such that the gauge field $A$ vanishes at the event horizon located at $r_{h}$
@@
The Smarr formula is traditionally derived either by applying Euler’s theorem for homogeneous functions to the "bulk" scaling of thermodynamic quantities [^1] Scaling arguments have been widely used in black hole thermodynamics to derive Smarr formulas, accommodating various matter contents and asymptotic behaviors (see, e.g., Lu ; Alfaro1 ). or by evaluating the Komar integral at spatial infinity and the horizon Kastor2 ; Kastor3 ; Tomas1 ; Tomas2 ; Tomas3 . In this work, we adopt a different perspective, interpreting black hole thermodynamics as the thermodynamics of the dual field theory in the strongly interacting regime, in accordance with the AdS/CFT correspondence. Consequently, we derive the Smarr formula using thermodynamic considerations and the Euclidean action Karch . To this end, let us consider the grand canonical potential, $\Omega$:
@@
In our context, $M$ represents the internal energy $U$ of the system. For what is coming it is important to note that, in principle, $\Omega$ is a function of the volume $V$, which, in the case of black holes in AdS spacetimes, corresponds to the spatial volume of the boundary metric where the dual field theory resides. For the AdSd Reissner-Nordström black hole, we compute the grand canonical potential $\Omega$ using the Euclidean on-shell action
@@
Equating the definition of $\Omega$ in (9) with the "on-shell value" of $\Omega$ (10), we obtain the Smarr formula for AdS black holes:
@@
In the last term of this expression, the radius of the event horizon $r_{h}$ should be understood as a function of the thermodynamic variables. In summary, the Smarr formula is an unusual relation among thermodynamic quantities.
Our objective is to develop a deeper physical understanding of the Smarr formula within the broader framework of the AdS/CFT correspondence. In particular, we are going to recast the Smarr formula (11) into a form that naturally generalizes the standard Euler relation, $U+pV=TS+\mu Q$, by incorporating an additional correction term arising from a geometric scale inherent to the system. As previously noted in Karch , this correction captures the fact that the dual field theory lives on a sphere rather than a flat plane. Our analysis will go beyond purely thermodynamic reasoning by incorporating insights from the fluid/gravity correspondence.
In thermodynamics, the Euler equation arises in the thermodynamic limit, where the volume $V$ and the number of particles $N$ both tend to infinity while maintaining a fixed ratio $\frac{N}{V}$. In this limit, the grand canonical potential becomes $\Omega=-pV$, and the Euler equation follows as $-pV=U-TS-\mu Q$. We seek to extend this reasoning to the Smarr formula. However, in the context of black hole thermodynamics, defining the thermodynamic volume $V$ and pressure $p$ is far from straightforward. This is where fluid/gravity duality offers crucial insight. To define the pressure $p$ for AdS black holes, we observe that the boundary stress-energy tensor associated with these black holes takes the form of a perfect fluid stress tensor Clifford0 ; Clifford . The boundary stress tensor for any given gravitational solution can be calculated using the formula provided in Kraus ; Clifford1 ; Skenderis :
$T_{ab}=\frac{1}{8\pi G_{N}}\left(h_{ab}\left(K-\frac{(d-2)}{l}\right)+K_{ab}+\frac{lG_{ab}}{(d-3)}+...\right)~{},$ (12)
where $h_{ab}$ is the induced metric, $K_{ab}$ is the extrinsic curvature, and $G_{ab}$ is the Einstein tensor for a fixed value of the radial direction $r$. The counterterms explicitly written in (12), are valid for $d<6$; for higher dimensions, see Clifford1 . As a reminder, $d$ is the bulk dimension, so $(d-1)$ is the boundary dimension. In our case, the boundary metric is
@@
Thus, the volume of the spatial section of the boundary spacetime is $V=l^{d-2}A_{d-2}$ where the AdS radius $l$ is acting as the radius of the $(d-2)$-dimensional sphere. For even-dimensional boundary spacetime, the conformal anomaly and Casimir energy appear Kraus ; Skenderis . Since these effects are not relevant to the present discussion, we restrict our analysis to odd-dimensional boundary spacetime. For $(d-1)=2n+1$ dimensions, the components of the boundary stress tensor are
$-T^{t}_{\;t}=\frac{(d-2)\tilde{m}}{16\pi G_{n}l^{d-2}}=\frac{M}{l^{d-2}A_{d-2}}=\frac{M}{V}~{},$ (14)
$T^{\phi_{j}}_{\;\phi_{j}}=\frac{\tilde{m}}{16\pi G_{n}l^{d-2}}=\frac{M}{(d-2)V}~{}.$ (15)
@@
$p=\frac{\epsilon}{(d-2)}~{},\quad\epsilon=\frac{M}{V}~{}.$ (17)
This holographic perfect fluid in certain sense is dual to the AdSd Reissner-Nordström black hole. Clearly, this fluid satisfies the conformal equation of state $\epsilon=(d-2)p$. The thermodynamics of this fluid is the same than the black hole thermodynamics since the fluid is at global thermal equilibrium.
@@
which resembles the Euler equation plus an additional correction term. As mentioned earlier, the traditional Euler equation arises in the thermodynamic limit, where $V\to\infty$. However, in our case, the volume $V$ occupied by the fluid is finite, and thus the standard derivation of the Euler equation no longer applies. In the gravitational context, the limit $V\to\infty$ corresponds to a charged black brane, for which the Smarr formula reduces to the Euler equation $p+\epsilon=Ts+\mu q$ Makoto , as the dual field theory lives in a plane.
In the analysis leading to equation (18), we relied on the idea that a perfect fluid in global thermal equilibrium provides an equivalent description of global thermodynamics. However, perfect fluids are typically used to describe local thermodynamics, where the Euler equation is assumed to hold universally—a point that seems at odds with our intuitive understanding of the Smarr formula developed in this section. To fully grasp the physical significance of the Smarr formula, it is essential to extend the Euler equation beyond its conventional domain. Therefore, in the next section, we will set aside the discussion of black holes and focus solely on the Euler equation and its generalization in the context of perfect fluids.
@@ 3 Effective action for perfect fluids
In this section, we begin by reviewing the standard construction of the effective action for perfect fluids, which naturally leads to the Euler equation. We then extend this framework by introducing a new variable into the effective action, capturing the fluid’s response to the presence of a geometrical scale. With this extension, we derive a generalized form of the Euler equation.
@@ 3.1 Review standard effective action for perfect fluids
Formal constructions of effective actions for hydrodynamics have been developed using techniques such as the Schwinger-Keldysh formalism Liu2 ; Liu3 ; Ranga1 ; Ranga2 and superspace methods Ranga3 ; Ranga4 ; Jensen1 . Comprehensive reviews can be found in Liu4 ; Basar . Although these advanced techniques will not be explicitly employed in this article, they highlight the significant progress and depth this field has achieved. Instead, we revisit the foundational constructions presented in Dubovsky1 ; Son ; Dubovsky2 , informed by a more modern understanding of the subject. Moreover, we focus exclusively on perfect fluids, leaving aside considerations of dissipation and hydrodynamic anomalies.
Hydrodynamics provides a valid effective description in the long-wavelength limit. For this reason, we assume that the curvature of the manifolds under consideration is much larger than the mean free path of the fluid. To construct the effective action for perfect fluids, two key components are required:
@@
Given the dynamical variables and symmetries of the effective action, we should, in principle, include all gauge-invariant terms in the theory. However, since our focus is on perfect fluids, we will disregard higher-derivative terms in the effective action. Moreover, as stated in the introduction, we assume that the bulk spacetime has d dimensions. Consequently, in this section, we consider a perfect fluid in an odd $(d-1)=2n+1$ dimensional manifold, ensuring that anomalies and Casimir energy do not play a role in our discussion. Finally, we limit this section to field-theoretical aspects, deferring the connection between these results and gravity to the next section.
Let’s begin by establishing the notation: the background metric for the perfect fluid is
@@
These symmetry transformations can be motivated by the observation that $\sigma^{I}$ represents spatial coordinates describing the fluids. As such, $\sigma^{I}$ must satisfy the comoving condition
@@
where $x^{a}$ in the position in the spacetime, and the fluid velocity is $u^{a}=\frac{dx^{a}}{d\tau}$. The second expression in (23) employs the directional derivative along the fluid velocity $u^{a}$. Notably, this condition is preserved by the symmetry transformations (21) and (22). It is also important to highlight that, unlike $\sigma^{I}$, the temporal scalar field satisfies $\frac{d}{d\tau}\sigma^{0}(x^{a}(\tau))\neq 0$, yet it remains gauge-invariant under these symmetry transformations.
In essence, the perfect fluid is characterized by a target metric $G_{MN}(\sigma^{M})$, where the fluid coordinates are $\sigma^{M}=(\sigma^{0},\sigma^{I})$. In this framework, the internal symmetries (21) and (22) correspond to the isometries of a degenerate metric Son :
$ds^{2}=G_{MN}d\sigma^{M}d\sigma^{N}~{},\quad G_{NM}n^{M}=0~{},\quad M,N\in\{0,1,...,d-2\}~{},$ (24)
@@
$ds^{2}=0\cdot(d\sigma^{0})^{2}+G_{IJ}(d\sigma^{I}-v^{I}d\sigma^{0})(d\sigma^{J}-v^{J}d\sigma^{0})~{},\quad I,J\in\{1,...,d-2\}~{},$ (25)
@@
Given the low-energy dynamical variables $\sigma^{M}$ and $\varphi$, along with the corresponding low-energy gauge symmetries, we proceed to construct the effective action for a charged perfect fluid. To achieve this, we must identify scalar quantities that remain invariant under diffeomorphisms of the physical metric $g_{ab}$ and are also gauge-invariant under the fluid symmetry transformations (21) and (22).
One such scalar is the entropy density $s$ of the system, which is defined in terms of the spatial target metric $G_{IJ}$ (25) as follows Son :
@@
Thus, the Euler equation is an unavoidable consequence of the effective action (32). Note that the equation of state for the fluid is not used at all in this derivation.
For completeness, we mention that due to the comoving condition (23), the entropy current $s^{a}$ is conserved off-shell: